OpenStudy (anonymous):

Got stuck on a integral, on which I can't seem to do partial fractions on. Integrate: 1 / (x (x^2+1)) Wolfram didn't really help me at all :(.

OpenStudy (anonymous):

Just go to show steps and it shows you what it does for partial fractions

OpenStudy (anonymous):

I don't understand how it did that. :/, For example it says u = x^2 and then the normal X becomes u?

OpenStudy (anonymous):

yup u substitution...have you done that before?

OpenStudy (anonymous):

Yes, I have but if you look at wolframs substition, u=x^2 but the normal x becomes u aswell.

OpenStudy (anonymous):

do you see the part that says du= 2x dx ???

OpenStudy (anonymous):

yes, does that allow x to become u?

OpenStudy (anonymous):

well isolate dx dx= (1/2x) du

OpenStudy (anonymous):

replace dx with that now

OpenStudy (anonymous):

and you already have a 1/x in there so 1/x * 1/x = 1/x^2

OpenStudy (anonymous):

and u=x^2

OpenStudy (anonymous):

so you get 1/u

OpenStudy (anonymous):

but you also see the 1/2 right that's where it came from

OpenStudy (anonymous):

HOLY pellet!, but the constant 1/2 goes outside the integral right?

OpenStudy (anonymous):

yup the half is put out

OpenStudy (anonymous):

wow thanks alot! understand it now!

OpenStudy (anonymous):

then the next part is partial fractions and I'll have to review it if you need help on that too. it's been a while.

OpenStudy (anonymous):

or w/e it called the integration by parts...

OpenStudy (anonymous):

partial fractions: A/u + B/u+1 and then find out A and B right?

OpenStudy (anonymous):

sounds about right....don't take my word for it

OpenStudy (anonymous):

:) thanks alot! for helping understand it now!.